### 04 Mar 2026 v1

# whose only singularity is a double point

Abstract.
We exhibit planar, rational curves of large degree over ${\mathbb{F}}_{2}$ that have a unique singular point, which has multiplicity 2. In characteristic 0 such curves exist only for degrees up to $6$.
The aim of this note is to construct rational curves $C_{d}\subset{\mathbb{F}}_{2}{\mathbb{P}}^{2}$ of degree $d=2^{n}+2$ that have a unique singular point, which has multiplicity 2; see Example 3.
In characteristic 0 such curves exist only for degrees up to $6$, see Lemma 4 and Remark 6.
For $d\geq 10$ the curves $C_{d}$ cannot be lifted to characteristic 0 in a way that preserves the standard resolution-of-singularities blow-up sequence; see Remark 7. This is inconsistent with my understanding of some steps of the program outlined in [Ish25]; see Remark 7. However, these curves do not seem to shed light on whether the set of log canonical thresholds depends on the characteristic; see Remark 8. This question is one of the main topics of [Ish25].
A general, planar, rational curve $B_{d}\subset{\mathbb{P}}^{2}$ of degree $d$ has $\binom{d-1}{2}$ nodes, and it is not easy to obtain examples with few singular points. There seem to be very few curves with only one singular point, which is a cusp. These appear prominently in the Abhyankar-Moh-Suzuki theorem describing $\operatorname{Aut}({\mathbb{C}}^{2})$ [Suz74, AM75]. Those with a single Puiseux pair are classified in [FdBLMHN07].
It is especially restrictive to have only cusps with multiplicity 2 [MS89, GZN00]; these can give surprising examples of real polynomials [JK11].
In positive characteristic, Artin-Schreier–type polynomials have been used to give non-standard embeddings of ${\mathbb{A}}^{1}$ into ${\mathbb{A}}^{2}$. This was most likely observed by Abhyankar; [Sat76] refers to them as ‘known counterexamples.’ The curves $C_{d}$ can be given by Artin-Schreier–type equations as in (2.5), but their properties are easier to see using an explicit parametrization.
Example 1.
Fix a prime $p$, let $q$ be a $p$-power and $r$ a multiple of $p$. Let $C^{\circ}_{q,r}\subset{\mathbb{A}}^{2}$ be the the image of the map
$\phi_{q,r}:t\mapsto\bigl(t^{q},t^{q+r}+t\bigr)=:(x,y),$
and $C_{q,r}\subset{\mathbb{P}}^{2}$ its closure.
Lemma 2.
For the curves $C_{q,r}$ in Example 1 the following hold.
  1. (2.1)
    $\phi_{q,r}:{\mathbb{A}}^{1}\to{\mathbb{A}}^{2}$ is a closed embedding,
  2. (2.2)
    $C^{\circ}_{q,r}$ is a smooth, rational curve, and
  3. (2.3)
    $C_{q,r}$ has a single point at infinity, which has multiplicity $r$.
Proof. The derivatve of $\phi_{q,r}$ is $(0,1)$, so it is an immersion. Since $t^{q}=x$, it is also purely inseparable, hence an isomorphism. The equation of $C^{\circ}_{q,r}$ is obtained using
(2.4) $y^{q}=t^{q(q+r)}+t^{q}=x^{q+r}+x.$
The unique point at infinity is (0:1:0)(0{:}1{:}0), and the equation there is
(2.5) $z^{r}=x^{q+r}+xz^{q+r-1}.\qed$
Example 3.
Setting $p=r=2,q=2^{n}$ we get
$C_{2^{n},2}:=\bigl(z^{2}y^{2^{n}}=x^{2^{n}+2}+xz^{2^{n}+1}\bigr)\subset{\mathbb{P}}^{2}.$
It is a rational curve of degree $2^{n}+2$. It has a unique singular point, which is a cusp of multiplicity 2.
It is thus of type $A_{m}$, where $m=2^{n}(2^{n}+1)$. The latter is roughly $(\deg C_{2^{n},2})^{2}$.
Note that for $m$ even, there are non-equivalent singularities of type $A_{m}$ in characteristic 2 with the same embedded resolution graphs. In the notation of [KS85, GK90], the normal form of the singularity of $C_{2^{n},2}$ is $z^{2}+zx^{r}+x^{m+1}$ where $m=2^{n}(2^{n}+1)$ and $r=\frac{1}{2}m+2^{n-1}+1$.
Lemma 4.
[GZN00] Let $B\subset{\mathbb{C}}{\mathbb{P}}^{2}$ be a reduced, (not necessarily rational) curve of degree $2d$ with a singularity of type $A_{m}$. Then $m\leq 3d(d-1)+1$, which is roughly $\frac{3}{4}(\deg B)^{2}$.
Hint of proof. Let $S\to{\mathbb{P}}^{2}$ be the double cover ramified along $B$. It has a surface singularity of type $A_{m}$. Thus the cohomology of the vanishing cycles of the $A_{m}$ surface singularity embeds into $H^{2,2}(S_{t},{\mathbb{C}})$ of a smoothing $S_{t}$ of $S$. ∎
Remark 5.
The upper bound in Lemma 4 is not expected to be sharp for $d>6$. The reducible curves
$D_{2d}:=\bigl((y-x^{d})^{2}=y^{2d}\bigr)$
have a singularity of type $A_{2d^{2}-1}$. This corresponds to $m=\frac{1}{2}(\deg D_{2d})^{2}-1$.
It is not easy to find curves $B$ with an $A_{m}$ singularity such that $m>\frac{1}{2}(\deg B)^{2}$. The largest known examples give asymptotically $m\approx\frac{15}{28}(\deg B)^{2}$ [GZN00].
Remark 6.
For $d\leq 6$ [Yos79] gives degree $d$, rational, plane curves $Y_{d}\subset{\mathbb{C}}{\mathbb{P}}^{2}$ with a single singular point, which has multiplicity 2.
For $d=6$ the equation is explicit, but rather complicated. The curve $Y_{6}$ has an $A_{19}$ singularity; the largest possible.
The curve $C_{4,2}$ has degree $6$ and an $A_{20}$ singularity. It is possible that $Y_{6}$ specializes to $C_{4,2}$. Note that the singularity type changes from $A_{19}$ to $A_{20}$, but they both need 10 blow-ups to resolve. So the embedded resolution of $C_{4,2}$ as in (7.1) might be liftable to an embedded resolution of $Y_{6}$.
Remark 7 (Connection with [Ish25]).
Let $0\in X\subset{\mathbb{A}}^{n}$ be a singularity. Assume for simplicity that it can be resolved by a sequence of blowing up points
(7.1) $(Y_{r},X_{r},p_{r})\to\cdots\to(Y_{1},X_{1},p_{1})\to(Y_{0},X_{0},p_{0})=({\mathbb{A}}^{n},X,0),$
where $X_{i}\subset Y_{i}$ is the birational transform of $X$, and $p_{i+1}$ lies over $p_{i}$. Let $\pi_{i}:Y_{i}\to Y_{0}$ denote the composite.
The discrepancies of the exceptional divisors with respect to the ideal sheaf of $X$ are determined by the numbers $\operatorname{mult}_{p_{i}}\pi_{i}^{-1}(p_{0})$ and $\operatorname{mult}_{p_{i}}X_{i}$.
Thus if we deform the blow-up sequence (7.1) such that these multiplicities stay constant, then the discrepancies of the $\pi_{r}$-exceptional divisors are invariant under this deformation. Call these discrepancy preserving deformations of (7.1).
[Ish25] notes that every exceptional divisor over ${\mathbb{A}}^{n}$ can be obtained as a $\pi_{r}$-exceptional divisor for a (somewhat more general) blow-up sequence as in (7.1).
This is especially useful if $0\in X\subset{\mathbb{A}}^{n}$ is in characteristic $p$, and the general fiber of such a deformation is in characteristic 0. It seems that in many cases there are discrepancy preserving liftings from characteristic $p$ to characteristic 0.
However, the curves in Example 3 show that this way of preserving discrepancies in liftings is not always possible.
To get such examples, let $C_{d}=\bigl(g(x,y,z)=0\bigr)$ be a curve of degree $d$ in characteristic $2$ that has a single $A_{2r}$ singularity. For odd $m\gg 1$ consider
$X_{g}:=\bigl(g(x,y,z)+x^{m}+y^{m}+z^{m}=0\bigr)\subset{\mathbb{A}}^{3}.$
This can be resolved by first blowing up the origin, and then $r$-times the singular point of the curve $C_{d}$.
The resulting blow-up sequence has a discrepancy preserving lifting to characteristic 0 iff the curve $C_{d}$ lifts to a curve $C^{\prime}_{d}$ in characteristic 0 that has an $A_{2r}$ or $A_{2r-1}$ singularity.
The smallest counterexample is $C_{8,2}$, which has degree $10$ and an $A_{72}$ singularity. By contrast, any reduced, decic curve $B_{10}\subset{\mathbb{C}}{\mathbb{P}}^{2}$ has at most an $A_{60}$ singularity by Lemma 4.
Remark 8 (Log canonical thresholds).
The log canonical threshold of the curve $C_{2^{n},2}\subset{\mathbb{F}}_{2}{\mathbb{P}}^{2}$ is $\frac{1}{2}+\frac{1}{m+1}$, where $m=2^{2n}+2^{n}$. For $n\geq 2$ there is no curve $B\subset{\mathbb{C}}{\mathbb{P}}^{2}$ with the same degree and same log canonical threshold as $C_{2^{n},2}$.
However, the log canonical threshold of the surface $X(g)\subset{\mathbb{A}}^{3}$ constructed from $C_{2^{n},2}$ is $3/(2^{n}+2)$. It is computed by the exceptional divisor of the first blow-up, and a general point of the same multiplicity has the same log canonical threshold.
Note, however, that these non-liftable examples seem rather exceptional, and possibly their discrepancies can be handled in different ways. The next examples show that liftability may frequently hold.
Example 9 (Quartic surfaces).
I first tried to look for non-liftable examples among K3 surfaces. In characteristic 0 the rank of the Néron-Severi group of a smooth K3 is at most 20, but in characteristic $p$ it is 22 for supersingular ones.
So if $S$ is a supersingular K3 surface with Picard number 1 in characteristic $p$, then there is no lifting to characteristic 0 preserving all the singularity types.
Most supersingular K3 surfaces with Picard number 1 have many singularities, and individually each singularity seems to lift, but there are a few examples where lifting even one singularity is not possible. I thank Shimada and Schütt for explaining these to me.
Working with the lists of [Shi04], we discuss 4 extremal cases, but none are good enough to be used in Remark 7.
(9.1) There is a degree 4 K3 surface $S_{1}$ in characteristic $2$ with a $D_{21}$ singularity. However, $S_{1}$ is not a quartic in ${\mathbb{P}}^{3}$, but a double cover of the quadric cone. Thus $S_{1}$ is naturally an octic surface in ${\mathbb{P}}(1,1,2,4)$. So there is an octic surface in ${\mathbb{P}}(1,1,2,4)$ with a $D_{21}$ singularity in characteristic $2$, but there is no such in char 0.
It seems likely that if $(s,S)$ is any quartic surface in characteristic $p$ with with an ADE singularity at $s$, then there is a lifting to characteristic 0 preserving the singularity type at $s$.
(9.2) There is a K3 surface $S_{2}$ in characteristic $11$ with an $A_{21}$ singularity, but it has degree $22$.
(9.3) There is a quasi-elliptic K3 surface $S_{3}$ in characteristic $2$ with a $D_{20}$ singularity, and a section along which the surface is smooth. The Weierstrass equation realizes it is a surface in the projectivization of ${\mathcal{O}}_{{\mathbb{P}}^{1}}+{\mathcal{O}}_{{\mathbb{P}}^{1}}(8)+{\mathcal{O}}_{{\mathbb{P}}^{1}}(12)$, but there is no similar surface in characteristic 0.
(9.4) There is an elliptic K3 surface $S_{4}$ in characteristic $2$ with $A_{17}+3A_{1}$ singularities, and a section along which the surface is smooth. As before, there is no lifting that keeps all singularities, but there is an elliptic K3 surface in characteristic 0 with $A_{17}+A_{1}$ singularities.
Acknowledgments.
I thank M. Schütt and I. Shimada for numerous e-mails explaining supersingular K3 surfaces. Partial financial support was provided by the Simons Foundation (grant number SFI-MPS-MOV-00006719-02).

### 16 Mar 2026 v2

@@
 In characteristic 0 such curves exist only for degrees up to $6$.
-The aim of this note is to construct rational curves $C_{d}\subset{\mathbb{F}}_{2}{\mathbb{P}}^{2}$ of degree $d=2^{n}+2$ that have a unique singular point, which has multiplicity 2; see Example 3.
+The aim of this note is to construct rational curves $C_{d}\subset{\mathbb{F}}_{2}{\mathbb{P}}^{2}$ of degree $d=2^{n}+2$ or $d=2^{n}$ that have a unique singular point, which has multiplicity 2; see Example 3.
 In characteristic 0 such curves exist only for degrees up to $6$, see Lemma 4 and Remark 6.
@@
 This is inconsistent with my understanding of some steps of the program outlined in [Ish25]; see Remark 7.
-However, these curves do not seem to shed light on whether the set of log canonical thresholds depends on the characteristic; see Remark 8.
-This question is one of the main topics of [Ish25].
+However, these curves do not seem to shed light on whether the set of log canonical thresholds depends on the characteristic, which is one of the main topics of [Ish25]; see Remark 8.
 A general, planar, rational curve $B_{d}\subset{\mathbb{P}}^{2}$ of degree $d$ has $\binom{d-1}{2}$ nodes, and it is not easy to obtain examples with few singular points.
@@
 Example 1.
-Fix a prime $p$, let $q$ be a $p$-power and $r$ a multiple of $p$. Let $C^{\circ}_{q,r}\subset{\mathbb{A}}^{2}$ be the the image of the map
-$\phi_{q,r}:t\mapsto\bigl(t^{q},t^{q+r}+t\bigr)=:(x,y),$
-and $C_{q,r}\subset{\mathbb{P}}^{2}$ its closure.
+Fix a characteristic $p>0$. Let $q$ be a $p$-power, and $g(u)\in{\mathbb{F}}_{p}[u]$ a polynomial. Let $C^{\circ}_{q,g}\subset{\mathbb{A}}^{2}$ be the the image of the map
+$\phi_{q,g}:t\mapsto\bigl(t^{q},g(t^{p})+t\bigr)=:(x,y),$
+and $C_{q,g}\subset{\mathbb{P}}^{2}$ its closure.
 Lemma 2.
-For the curves $C_{q,r}$ in Example 1 the following hold.
+For the curves $C_{q,g}$ in Example 1 the following hold.
   1. (2.1)
-    $\phi_{q,r}:{\mathbb{A}}^{1}\to{\mathbb{A}}^{2}$ is a closed embedding,
+    $\phi_{q,g}:{\mathbb{A}}^{1}\to{\mathbb{A}}^{2}$ is a closed embedding,
   2. (2.2)
-    $C^{\circ}_{q,r}$ is a smooth, rational curve, and
+    $C^{\circ}_{q,g}$ is a smooth, rational curve, and
   3. (2.3)
-    $C_{q,r}$ has a single point at infinity, which has multiplicity $r$.
-Proof. The derivatve of $\phi_{q,r}$ is $(0,1)$, so it is an immersion. Since $t^{q}=x$, it is also purely inseparable, hence an isomorphism. The equation of $C^{\circ}_{q,r}$ is obtained using
-(2.4) $y^{q}=t^{q(q+r)}+t^{q}=x^{q+r}+x.$
-The unique point at infinity is (0:1:0)(0{:}1{:}0), and the equation there is
-(2.5) $z^{r}=x^{q+r}+xz^{q+r-1}.\qed$
+    $C_{q,g}$ has a single point at infinity. If $r:=p\cdot\deg g-q$ is positive, it has multiplicity $r$.
+Proof. The derivatve of $\phi_{q,g}$ is $(0,1)$, so it is an immersion. Since $t^{q}=x$, it is also purely inseparable, hence an isomorphism. The equation of $C^{\circ}_{q,g}$ is obtained using
+(2.4) $y^{q}=g(t^{qp})+t^{q}=g(x^{p})+x.$
+If $r:=p\cdot\deg g-q$ is positive, then the unique point at infinity is (0:1:0)(0{:}1{:}0), and the equation there is
+(2.5) $z^{r}=g(x^{p})+xz^{q+r-1}.\qed$
 Example 3.
-Setting $p=r=2,q=2^{n}$ we get
+Setting $p=2,q=2^{n}$ and $g(u)=u^{2^{n-1}+1}$, we get
 $C_{2^{n},2}:=\bigl(z^{2}y^{2^{n}}=x^{2^{n}+2}+xz^{2^{n}+1}\bigr)\subset{\mathbb{P}}^{2}.$
@@
 In the notation of [KS85, GK90], the normal form of the singularity of $C_{2^{n},2}$ is $z^{2}+zx^{r}+x^{m+1}$ where $m=2^{n}(2^{n}+1)$ and $r=\frac{1}{2}m+2^{n-1}+1$.
+Setting $p=2,q=2^{n}$ and $g(u)=u^{2^{n-1}-1}$, we get
+$C_{2^{n}}:=\bigl(y^{2^{n}}=z^{2}+z^{2^{n}-1}\bigr)\subset{\mathbb{P}}^{2}.$
+It is a rational curve of degree $2^{n}$. It has a unique singular point of type $A_{m}$, where $m=(2^{n}-1)(2^{n}-2)$.
 Lemma 4.
@@
 have a singularity of type $A_{2d^{2}-1}$. This corresponds to $m=\frac{1}{2}(\deg D_{2d})^{2}-1$.
-It is not easy to find curves $B$ with an $A_{m}$ singularity such that $m>\frac{1}{2}(\deg B)^{2}$. The largest known examples give asymptotically $m\approx\frac{15}{28}(\deg B)^{2}$ [GZN00].
+It is not easy to find curves $B$ with an $A_{m}$ singularity such that $m>\frac{1}{2}(\deg B)^{2}$. The largest known examples give asymptotically $m\approx\frac{7}{12}(\deg B)^{2}$ [Ore12, Sec.5].
 Remark 6.
 For $d\leq 6$ [Yos79] gives degree $d$, rational, plane curves $Y_{d}\subset{\mathbb{C}}{\mathbb{P}}^{2}$ with a single singular point, which has multiplicity 2.
-For $d=6$ the equation is explicit, but rather complicated. The curve $Y_{6}$ has an $A_{19}$ singularity; the largest possible.
+The curve $Y_{6}$ has an $A_{19}$ singularity; the largest possible.
+Simpler equations are given in [ABCRCA02, 5.8].
 The curve $C_{4,2}$ has degree $6$ and an $A_{20}$ singularity.
@@
 The resulting blow-up sequence has a discrepancy preserving lifting to characteristic 0 iff the curve $C_{d}$ lifts to a curve $C^{\prime}_{d}$ in characteristic 0 that has an $A_{2r}$ or $A_{2r-1}$ singularity.
-The smallest counterexample is $C_{8,2}$, which has degree $10$ and an $A_{72}$ singularity.
-By contrast, any reduced, decic curve $B_{10}\subset{\mathbb{C}}{\mathbb{P}}^{2}$ has at most an $A_{60}$ singularity by Lemma 4.
+The smallest counterexample is $C_{8}$, which has degree $8$ and an $A_{42}$ singularity.
+A reduced, octic curve $B_{8}\subset{\mathbb{C}}{\mathbb{P}}^{2}$ has at most an $A_{37}$ singularity by Lemma 4.
+The next one is $C_{8,2}$, which has degree $10$ and an $A_{72}$ singularity. A reduced, decic curve $B_{10}\subset{\mathbb{C}}{\mathbb{P}}^{2}$ has at most an $A_{61}$ singularity by Lemma 4.
 Remark 8 (Log canonical thresholds).
+The program of Ishii aims to understand log canonical thresholds in positive characteristic [Ish25].
 The log canonical threshold of the curve $C_{2^{n},2}\subset{\mathbb{F}}_{2}{\mathbb{P}}^{2}$ is $\frac{1}{2}+\frac{1}{m+1}$, where $m=2^{2n}+2^{n}$.
@@
 However, the log canonical threshold of the surface $X(g)\subset{\mathbb{A}}^{3}$ constructed from $C_{2^{n},2}$ is $3/(2^{n}+2)$.
-It is computed by the exceptional divisor of the first blow-up, and a general point of the same multiplicity has the same log canonical threshold.
-Note, however, that these non-liftable examples seem rather exceptional, and possibly their discrepancies can be handled in different ways. The next examples show that liftability may frequently hold.
+It is computed by the exceptional divisor of the first blow-up, and a general singular point with the same multiplicity has the same log canonical threshold.
+Thus these curves do not seem to give further information about the program in [Ish25].
+Note, however, that the above non-liftable examples seem rather exceptional, and possibly their discrepancies can be handled in different ways.
+The next examples show that liftability may frequently hold.
 Example 9 (Quartic surfaces).
@@
 As before, there is no lifting that keeps all singularities, but there is an elliptic K3 surface in characteristic 0 with $A_{17}+A_{1}$ singularities.
+The curves $C_{2^{n},2}$ also lead to supersingular double planes.
+10Supersingular double planes.
+Over ${\mathbb{F}}_{2}$ consider the surfaces
+(10.1) $S_{r}:=\bigl(u^{2}+u(zy^{2r}+x^{2r+1})+xz^{4r+1}+z^{2}y^{4r}+x^{4r+2}=0\bigr)\subset{\mathbb{P}}(1,1,1,2r+1).$
+In the $y=1$ chart, the equation is
+(10.2) $g:=u^{2}+u(z+x^{2r+1})+xz^{4r+1}+z^{2}+x^{4r+2}=0.$
+Thus $g_{u}=z+x^{2r+1},\ g_{z}=u+xz^{4r},\ g_{x}=ux^{2r}+z^{4r+1}.$ The first 2 define the monomial curve $M_{r}$ which is the image of
+$\phi_{r}:t\mapsto\bigl(t,t^{2r+1},t^{8r^{2}+4r+1}\bigr),$
+and $g_{x}$ vanishes identically along $M_{r}$. Thus
+$\begin{array}[]{rcl}{\mathbb{F}}_{2}[x,z,u]/(g,g_{x},g_{z},g_{u})&\cong&{\mathbb{F}}_{2}[t]/\bigl(g(t,t^{2r+1},t^{8r^{2}+4r+1})\bigr)\\ &\cong&{\mathbb{F}}_{2}[t]/\bigl(t^{8r^{2}+6r+2}(t^{8r^{2}+2r}+1)\bigr).\end{array}$
+Thus $S_{r}$ has an $A_{8r^{2}+6r+1}$ singularity at $t=0$.
+For the other singularities, write $2r=qr^{\prime}$ where $q$ is a $2$-power and $r^{\prime}$ is odd.
+Then we have $A_{q-1}$ singularities when $t$ is a $(4rr^{\prime}+r^{\prime})^{th}$ root of unity.
+There are no other singular points.
+If $r$ is a $2$-power, then the minimal resolution $\bar{S}_{r}\to S_{r}$ contains $8r^{2}+6r+1+(4r+1)(2r-1)=16r^{2}+4r$ exceptional curves.
+Over the line $(z=0)$ the equation (10.2) becomes $u^{2}+ux^{2r+1}+x^{4r+2}=0$.
+Setting $v:=ux^{-2r-1}$ we get $v^{2}+v+1=0$.
+If $\alpha_{1},\alpha_{2}$ are its roots, then the preimage of $(z=0)$ has 2 components
+$B_{i}:=\bigl(u+\alpha_{i}x^{2r+1}=z=0\bigr).$
+Thus the Picard number of $\bar{S}_{r}$ is at least $16r^{2}+4r+2$.
+This equals the second Betti bumber of a double plane ramified along a smooth curve of degree $4r+2$.
+Thus the surfaces $\bar{S}_{r}$ are supersingular for every $r=2^{n}$, with Picard number $16r^{2}+4r+2$.
+Example 11.
+The surface $S_{1}$ is a K3 surface with $A_{15}+5A_{1}$ singularities.
+Resolving the $A_{15}$ singularity, the birational transforms of the $B_{i}$ become smooth rational curves, and the dual configuration is
+$\begin{array}[]{ccccccccccccc}&&&&\bar{B}_{1}&&&&\bar{B}_{2}&&&&\\ &&&&|&&&&|&&&&\\ C_{1}&-&C_{2}&-&C_{3}&-&\cdots&-&C_{13}&-&C_{14}&-&C_{15}\end{array}$
+From this we compute that $(B_{1}\cdot B_{2})=\frac{9}{16}$ and $(B_{1}^{2})=(B_{2}^{2})=\frac{7}{16}$.
+Contracting all but the curve $C_{8}$ gives a supersingular K3 surface $S^{\prime}_{1}$ with singularities $2E_{8}+5A_{1}$, and Picard number $1$.
+The image of $C_{8}$ is a rational curve $C^{\prime}_{8}$ with 2 cusps at the $E_{8}$-points, and $(C^{\prime}_{8}\cdot C^{\prime}_{8})=2$.
+Thus $S^{\prime}_{1}$ is a K3 surface of degree 2.
+This is the 4th surface on the list of [Shi04] in characteristic 2.
 Acknowledgments.
-I thank M. Schütt and I. Shimada for numerous e-mails explaining supersingular K3 surfaces. Partial financial support was provided by the Simons Foundation (grant number SFI-MPS-MOV-00006719-02).
+I thank M.
+Schütt and I.
+Shimada for numerous e-mails explaining supersingular K3 surfaces, and S.
+Orevkov and M.
+Zaidenberg for references.
+Partial financial support was provided by the Simons Foundation (grant number SFI-MPS-MOV-00006719-02).